OpenStudy (anonymous):

WILL MEDAL! Suppose p varies directly with d, and p=7 when d=3 What is the value of d, when p=28

OpenStudy (karon86):

I'm here

OpenStudy (karon86):

Use cross multiplication

OpenStudy (anonymous):

?

OpenStudy (anonymous):

show me how

OpenStudy (karon86):

The formula for this is y1/x1=y2/x2

OpenStudy (karon86):

y=p x=d

OpenStudy (karon86):

So just plug the values in the variables. 7/3=28/x

OpenStudy (anonymous):

Oh ok

OpenStudy (karon86):

Use cross multiplication to solve for x. 3 x 28 = 84

OpenStudy (karon86):

Then just divide 84 by 7

OpenStudy (anonymous):

12

OpenStudy (karon86):

x = 12 :)

OpenStudy (anonymous):

thanks man

OpenStudy (karon86):

Goodjob :D You welcome :D

OpenStudy (karon86):

Thanks :D

OpenStudy (anonymous):

ur welcome